OpenStudy (anonymous):

If y+2q=15, q+2p=5, p+2y=7, Then find the value of p+q+y. Please help! (:

OpenStudy (anonymous):

Just add all of them, you'll end up with an answer.

OpenStudy (anonymous):

We tried multiple times, but if you do that you will still have variables, and thats why it doesnt work.

OpenStudy (anonymous):

If you understand how to do it the way you are talking about, can you please explain it?

OpenStudy (anonymous):

Let me show you: y+2q+q+2p+p+2y=15+5+7 3y+3p+3q=22 3(y+p+q)=22

OpenStudy (anonymous):

How do you find the value of each variable then?

OpenStudy (anonymous):

We don't have to. What asked is p+q+y, and we already know that p+q+y=22/3

OpenStudy (anonymous):

We checked the answer, and that is not the answer. It is 9. You have to find the value of each variable, and add them to get 9. We checked the answer because we couldn't understand it

OpenStudy (anonymous):

Easiest way to find the answers is to create a matrix

OpenStudy (anonymous):

Once you make/solve a matrix you'll find that (x,y,z) = (-1,7,4)

jimthompson5910 (jim_thompson5910):

it should be 3(y+p+q)=27, not 3(y+p+q)=22

jimthompson5910 (jim_thompson5910):

but imron07 is on the right track

OpenStudy (anonymous):

answer should be 10 unless my math/copying skills suck